OpenStudy (anonymous):

Multiply and simplify: (7 + 3√2)(9 - 5√2)

OpenStudy (anonymous):

\[63-35\sqrt{2}+27\sqrt{2}-(15\times{2})\]\[63-8\sqrt{2}-30\]\[33-8\sqrt{2}\]

OpenStudy (anonymous):

you use the distributive property and multiply then you add the radical coefficients lastly we add the constants

OpenStudy (anonymous):

@SkyeJ helpful?? :)

OpenStudy (anonymous):

very helpful!!! thank you so much :D